How much rust (Fe 2 O 3 ) will be formed, when 100 kg of iron have completely rusted away? ( Atomic weights: Fe = 56 ; O = 16; H = 1)

Grade:12th pass

1 Answers

Sunil Kumar FP
9 years ago
rust (Fe 2 O 3 ) will be formed, when 100 kg of iron have completely rusted away
we have
2Fe + 3/2O2 ----Fe2O3
mole of iron in 100 kg of iron =100*10^3/56
mole of Fe2O3 formed =1/2 (mole of iron)
=(10^5)/56*2
weight of the iron oxide= (159.69 *10^5)/112
=1.425*10^5 gm